# Upper Approximation Bounds for Neural Oscillators

### v1

@@ A.2 Proof of Lemma 2
Proof: For every positive integer $M$ larger than one and arbitrary $M$ real frequencies $\omega_{1}$, $\omega_{2}$,…, $\omega_{M}$, with $\sigma_{\mathrm{ReLU}}(x)-\sigma_{\mathrm{ReLU}}(-x)=x$, there exists a set of equations with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$
@@ A.3 Proof of Lemma 3
Proof: From Lemma 1 and Lemma 2, for an arbitrary positive $v<T$, and every positive integer $M$ larger than one, there exists an MLP ${\it{\Gamma}}(\cdot)$ described in Lemma 2, such that for every $\mathbf{u}(t)\in K$, its corresponding solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from Eq. (1) described in Lemma 2 satisfies
@@
Since $\Phi$ is a causal and uniformly continuous operator from $C_{0}([0,T];{\mathbb{R}^{p}})$ to $C_{0}([0,T];{\mathbb{R}^{q}})$, and the parameters $\omega_{n}$, $\alpha_{n}$, and $\theta_{n}$ in Eq. (54) are independent of the input function $\mathbf{u}(t)$ according to the proof of Lemma 1, ${\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$ can be regarded as a continuous function with respect to the values of $\mathbf{x}(t)$, $\mathbf{u}(0)$, and time $t$. The continuity of ${\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$ with respect to the values of $\mathbf{x}(t)$ and $\mathbf{u}(0)$ is evident because $\Phi$ is a continuous operator. The continuity with respect to the value of $t$ is analytically established through Eq. (74) in the proof of Lemma 4 in A.4. Thus, for every positive integer M larger than one, a related continuous mapping $\Psi:{\mathbb{R}}^{p(M+1)}\times{[0,T]}\to{\mathbb{R}}^{q}$ can be defined as $\Psi(\mathbf{x},\mathbf{u}_{0},t)=\Psi[\mathbf{x}(t),\mathbf{u}(0),t]={\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$, where $\mathbf{x}=[\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top},...,\mathbf{x}_{M}^{\top}]^{\top}$, $\mathbf{x}_{n}$ and $\mathbf{u}_{0}$ represent the values of $\mathbf{x}_{n}(t)$ and $\mathbf{u}(0)$ in Eq. (54), respectively.
@@
$\begin{array}[]{l}|{\Phi}[\mathbf{u}(\tau)](t_{0})-\Psi(\mathbf{x},\mathbf{u}_{0},t_{0})|=|{\Phi}[\mathbf{u}(\tau)](t_{0})-{\Phi}[{\hat{\mathbf{u}}}_{t_{0}}(\tau)](t_{0})|\leq{\left\lVert{\Phi}[\mathbf{u}(\tau)](t)-{\Phi}[{\hat{\mathbf{u}}}_{t_{0}}(\tau)](t)\right\rVert}_{L^{\infty}[0,t_{0}]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}{\phi_{\Phi}}\left[{\left\lVert\mathbf{u}(t)-\hat{\mathbf{u}}_{t_{0}}(t)\right\rVert}_{L^{\infty}[0,t_{0}]}\right]\mathop{\leq}\limits^{(b)}{\phi_{\Phi}}(\varepsilon_{K}),\end{array}$ (56)
@@ A.4 Proof of Lemma 4
where $x_{n,i}(t)$ and $u_{i}(0)$ are respectively the $i^{\text{th}}$ elements of $\mathbf{x}_{n}(t)$ and $\mathbf{u}(0)$. Consider two approximates $\hat{\mathbf{u}}_{t,1}(\tau)$} and $\hat{\mathbf{u}}_{t,2}(\tau)$} of ${\mathbf{u}}_{1}(t)$ and ${\mathbf{u}}_{2}(t)$ $\in K$, respectively, with $\mathbf{x}_{1,n}(t)=\mathcal{L}_{t}\mathbf{u}_{1}(\omega_{n})$ and $\mathbf{x}_{2,n}(t)=\mathcal{L}_{t}\mathbf{u}_{2}(\omega_{n})$, from Eq. (58), the difference between the $i^{\text{th}}$ elements ${\hat{u}_{t,1,i}}(\tau)$ of $\hat{\mathbf{u}}_{t,1}(\tau)$} and ${\hat{u}_{t,2,i}}(\tau)$ of $\hat{\mathbf{u}}_{t,2}(\tau)$} can be bounded by
$\begin{array}[]{l}\mathop{\sup}\limits_{\tau\in[0,t]}\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|\\ \;\;\leq\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|\\ \;\;+\mathop{\sup}\limits_{\tau\in[0,t]}\left({\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}\left\{{\left[{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right]+\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|}}\right)\\ \;\;\leq\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left[{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right]+\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right|}\\ \;\;\mathop{\leq}\limits^{(a)}\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\frac{1}{T}\sum\limits_{n=1}^{M}{\left[{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|+2\left|{\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}}\right|}\right]}\\ \;\;\leq\left({1+\frac{{2M}}{{T{\omega_{1}}}}}\right)\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\frac{1}{T}\sum\limits_{n=1}^{M}{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|}\\ \;\;\mathop{\leq}\limits^{(b)}\max\left[{\left({1+2M}\right),{T^{-1}}}\right]\left[{\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\sum\limits_{n=1}^{M}{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|}}\right],\end{array}$ (59)
@@
$\begin{array}[]{l}\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\\ \;\;\mathop{\leq}\limits^{(a)}{\phi_{{\Phi_{j}}}}\left\{{\left\lVert{\hat{\mathbf{u}}_{t,1}}(\tau)-{\hat{\mathbf{u}}_{t,2}}(\tau)\right\rVert}_{L^{\infty}[0,t]}\right\}={\phi_{{\Phi_{j}}}}\left({\mathop{\sup}\limits_{\tau\in[0,t]}\left[{\sum\limits_{i=1}^{p}{\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|}}\right]}\right)\\ \;\;\leq{\phi_{{\Phi_{j}}}}\left[{\sum\limits_{i=1}^{p}{\mathop{\sup}\limits_{\tau\in[0,t]}\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|}}\right]\leq{\phi_{{\Phi_{j}}}}\left\{{\max\left[{\left({1+2M}\right),{T^{-1}}}\right]\left({\left|{\Delta{\bf{x}}}\right|+\left|{\Delta{{\bf{u}}_{0}}}\right|}\right)}\right\},\end{array}$ (60)
@@
$\begin{array}[]{l}\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\\ \;\;\;\leq\left|{{\Phi_{j}}\left[{{{{\hat{\bf u}}}_{t+\Delta t}}(\tau)}\right](t+\Delta t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)}\right|+\left|{{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right](t)}\right|\\ \;\;\;\leq\left|{\Delta t}\right|{D_{\Phi_{j}}}+\left|{{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right](t)}\right|\\ \;\;\;\leq\left|{\Delta t}\right|{D_{\Phi_{j}}}+{\phi_{{\Phi_{j}}}}\left[{{{\left\|{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)-{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right\|}_{{L^{\infty}}\left[0,t\right]}}}\right],\end{array}$ (64)
@@
Based Eqs. (62) and (63), the difference between ${\hat{\mathbf{u}}}_{t}(\tau)$ and ${{\hat{\mathbf{u}}}_{t+\Delta t}}(\tau)$, which is in the last term of Eq. (64), can be calculated as
$\begin{array}[]{l}{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)-{{{\hat{\mathbf{u}}}}_{t}}(\tau)=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\left\{{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]-1}\right\}}\right)\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\sum\limits_{n=1}^{M}{{\alpha_{n}}\left\{{{{\bf{x}}_{n}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\cos\left[{{\omega_{n}}(t+\Delta t)}\right]}\right)\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\sum\limits_{n=1}^{M}{{\alpha_{n}}\left[{{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\cos({\omega_{n}}t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right){A_{n}}(\Delta t)}+\sum\limits_{n=1}^{M}{{\alpha_{n}}\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}{B_{n}}(\Delta t)},\end{array}$ (65)
@@
$\begin{array}[]{l}\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right)}\right|\left|{{A_{n}}(\Delta t)}\right|}\\ \;\;\;\leq\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right)}\right|\left|{{\omega_{n}}\Delta t}\right|}\\ \;\;\;\leq\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{\omega_{n}}{{\bf{x}}_{n}}-{\alpha_{n}}{{\bf{u}}_{0}}}\right|}\\ \;\;\;\leq\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{\omega_{n}}{{\bf{x}}_{n}}}\right|}+\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{{\bf{u}}_{0}}}\right|}\\ \;\;\;\leq\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{{\omega_{n}}}\right|\left|{{{\bf{x}}_{n}}}\right|}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}\\ \;\;\;\leq\left|{\Delta t}\right|pT{B_{K}}\sum\limits_{n=1}^{+\infty}{\left|{{\alpha_{n}}}\right|\left|{{\omega_{n}}}\right|}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}\\ \;\;\;\mathop{\leq}\limits^{(a)}\left|{\Delta t}\right|pT{B_{K}}\frac{{{\pi^{2}}M{c_{\rho,3}}}}{{24T^{2}{c_{\rho,2}}}}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}=\left|{\Delta t}\right|p{B_{K}}MT^{-1}\left({\frac{{{\pi^{2}}{c_{\rho,3}}}}{{24{c_{\rho,2}}}}+1}\right),\end{array}$ (69)

### v2

@@ A.2 Proof of Lemma 2
Proof: For every positive integer $M$ larger than two and arbitrary $M$ real frequencies $\omega_{1}$, $\omega_{2}$,…, $\omega_{M}$, with $\sigma_{\mathrm{ReLU}}(x)-\sigma_{\mathrm{ReLU}}(-x)=x$, there exists a set of equations with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$
@@ A.3 Proof of Lemma 3
Proof: From Lemma 1 and Lemma 2, for an arbitrary positive $v<T$, and every positive integer $M$ larger than two, there exists an MLP ${\it{\Gamma}}(\cdot)$ described in Lemma 2, such that for every $\mathbf{u}(t)\in K$, its corresponding solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from Eq. (1) described in Lemma 2 satisfies
@@
Since $\Phi$ is a causal and uniformly continuous operator from $C_{0}([0,T];{\mathbb{R}^{p}})$ to $C_{0}([0,T];{\mathbb{R}^{q}})$, and the parameters $\omega_{n}$, $\alpha_{n}$, and $\theta_{n}$ in Eq. (59) are independent of the input function $\mathbf{u}(t)$ according to the proof of Lemma 1, ${\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$ can be regarded as a continuous function with respect to the values of $\mathbf{x}(t)$, $\mathbf{u}(0)$, and time $t$. The continuity of ${\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$ with respect to the values of $\mathbf{x}(t)$ and $\mathbf{u}(0)$ is evident because $\Phi$ is a continuous operator. The continuity with respect to the value of $t$ is analytically established through Eq. (79) in the proof of Lemma 4 in A.4. Thus, for every positive integer M larger than two, a related continuous mapping $\Psi:{\mathbb{R}}^{p(M+1)}\times{[0,T]}\to{\mathbb{R}}^{q}$ can be defined as $\Psi(\mathbf{x},\mathbf{u}_{0},t)=\Psi[\mathbf{x}(t),\mathbf{u}(0),t]={\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$, where $\mathbf{x}=[\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top},...,\mathbf{x}_{M}^{\top}]^{\top}$, $\mathbf{x}_{n}$ and $\mathbf{u}_{0}$ represent the values of $\mathbf{x}_{n}(t)$ and $\mathbf{u}(0)$ in Eq. (59), respectively.
@@
$|{\Phi}[\mathbf{u}(\tau)](t_{0})-\Psi(\mathbf{x},\mathbf{u}_{0},t_{0})|=|{\Phi}[\mathbf{u}(\tau)](t_{0})-{\Phi}[{\hat{\mathbf{u}}}_{t_{0}}(\tau)](t_{0})|\leq{\left\lVert{\Phi}[\mathbf{u}(\tau)](t)-{\Phi}[{\hat{\mathbf{u}}}_{t_{0}}(\tau)](t)\right\rVert}_{L^{\infty}[0,t_{0}]}$ (61)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}{\phi_{\Phi}}\left[{\left\lVert\mathbf{u}(t)-\hat{\mathbf{u}}_{t_{0}}(t)\right\rVert}_{L^{\infty}[0,t_{0}]}\right]\mathop{\leq}\limits^{(b)}{\phi_{\Phi}}(\varepsilon_{K}),$
@@ A.4 Proof of Lemma 4
where $x_{n,i}(t)$ and $u_{i}(0)$ are respectively the $i^{\text{th}}$ elements of $\mathbf{x}_{n}(t)$ and $\mathbf{u}(0)$. Consider two approximations $\hat{\mathbf{u}}_{t,1}(\tau)$} and $\hat{\mathbf{u}}_{t,2}(\tau)$} of ${\mathbf{u}}_{1}(t)$ and ${\mathbf{u}}_{2}(t)$ $\in K$, respectively, with $\mathbf{x}_{1,n}(t)=\mathcal{L}_{t}\mathbf{u}_{1}(\omega_{n})$ and $\mathbf{x}_{2,n}(t)=\mathcal{L}_{t}\mathbf{u}_{2}(\omega_{n})$, from Eq. (63), the difference between the $i^{\text{th}}$ elements ${\hat{u}_{t,1,i}}(\tau)$ of $\hat{\mathbf{u}}_{t,1}(\tau)$} and ${\hat{u}_{t,2,i}}(\tau)$ of $\hat{\mathbf{u}}_{t,2}(\tau)$} can be bounded by
$\mathop{\sup}\limits_{\tau\in[0,t]}\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|$ (64)
$\;\;\;\;\leq\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\mathop{\sup}\limits_{\tau\in[0,t]}\left({\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}\left\{{\left[{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right]+\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|}}\right)$
$\;\;\;\;\leq\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left[{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right]+\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right|}$
$\;\;\;\;\mathop{\leq}\limits^{(a)}\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\frac{1}{T}\sum\limits_{n=1}^{M}{\left[{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|+2\left|{\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}}\right|}\right]}$
$\;\;\;\;\leq\left({1+\frac{{2M}}{{T{\omega_{1}}}}}\right)\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\frac{1}{T}\sum\limits_{n=1}^{M}{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|}$
$\;\;\;\;\mathop{\leq}\limits^{(b)}\max\left[{\left({1+2M}\right),{T^{-1}}}\right]\left[{\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\sum\limits_{n=1}^{M}{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|}}\right],$
@@
$\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|$ (65)
$\;\;\;\;\;\mathop{\leq}\limits^{(a)}{\phi_{{\Phi_{j}}}}\left\{{\left\lVert{\hat{\mathbf{u}}_{t,1}}(\tau)-{\hat{\mathbf{u}}_{t,2}}(\tau)\right\rVert}_{L^{\infty}[0,t]}\right\}={\phi_{{\Phi_{j}}}}\left({\mathop{\sup}\limits_{\tau\in[0,t]}\left[{\sum\limits_{i=1}^{p}{\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|}}\right]}\right)$
$\;\;\;\;\;\leq{\phi_{{\Phi_{j}}}}\left[{\sum\limits_{i=1}^{p}{\mathop{\sup}\limits_{\tau\in[0,t]}\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|}}\right]\leq{\phi_{{\Phi_{j}}}}\left\{{\max\left[{\left({1+2M}\right),{T^{-1}}}\right]\left({\left|{\Delta{\bf{x}}}\right|+\left|{\Delta{{\bf{u}}_{0}}}\right|}\right)}\right\},$
@@
$\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|$ (69)
$\;\;\;\leq\left|{{\Phi_{j}}\left[{{{{\hat{\bf u}}}_{t+\Delta t}}(\tau)}\right](t+\Delta t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)}\right|+\left|{{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right](t)}\right|$
$\;\;\;\leq\left|{\Delta t}\right|{D_{\Phi_{j}}}+\left|{{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right](t)}\right|$
$\;\;\;\leq\left|{\Delta t}\right|{D_{\Phi_{j}}}+{\phi_{{\Phi_{j}}}}\left[{{{\left\|{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)-{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right\|}_{{L^{\infty}}\left[0,t\right]}}}\right],$
@@
From Eqs. (67) and (68), the difference between ${\hat{\mathbf{u}}}_{t}(\tau)$ and ${{\hat{\mathbf{u}}}_{t+\Delta t}}(\tau)$, which is in the last term of Eq. (69), can be calculated as
${{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)-{{{\hat{\mathbf{u}}}}_{t}}(\tau)=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\left\{{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]-1}\right\}}\right)\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]}$ (70)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\sum\limits_{n=1}^{M}{{\alpha_{n}}\left\{{{{\bf{x}}_{n}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\cos\left[{{\omega_{n}}(t+\Delta t)}\right]}\right)\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\sum\limits_{n=1}^{M}{{\alpha_{n}}\left[{{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\cos({\omega_{n}}t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right){A_{n}}(\Delta t)}+\sum\limits_{n=1}^{M}{{\alpha_{n}}\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}{B_{n}}(\Delta t)},$
@@
$\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right)}\right|\left|{{A_{n}}(\Delta t)}\right|}$ (74)
$\;\;\;\leq\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right)}\right|\left|{{\omega_{n}}\Delta t}\right|}=\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{\omega_{n}}{{\bf{x}}_{n}}-{\alpha_{n}}{{\bf{u}}_{0}}}\right|}\leq\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{\omega_{n}}{{\bf{x}}_{n}}}\right|}+\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{{\bf{u}}_{0}}}\right|}$
$\;\;\;\leq\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{{\omega_{n}}}\right|\left|{{{\bf{x}}_{n}}}\right|}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}\leq\left|{\Delta t}\right|pT{B_{K}}\sum\limits_{n=1}^{+\infty}{\left|{{\alpha_{n}}}\right|\left|{{\omega_{n}}}\right|}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}$
$\;\;\;\mathop{\leq}\limits^{(a)}\left|{\Delta t}\right|pT{B_{K}}\frac{{2c_{\rho,3}M^{3}}}{3T^{2}}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}\leq\left|{\Delta t}\right|p{B_{K}}M^{3}T^{-1}\left({c_{\rho,3}+1}\right),$